# Bob’s Confetti: Phonetic Memorization Attacks in Music and Video Generation

### v3

@@ Abstract
Generative AI systems for music and video commonly use text-based filters to prevent the regurgitation of copyrighted material. We expose a fundamental flaw in this approach by introducing Adversarial PhoneTic Prompting (APT), a novel attack that bypasses these safeguards by exploiting phonetic memorization. The APT attack replaces iconic lyrics with homophonic but semantically unrelated alternatives (e.g., “mom’s spaghetti” becomes “Bob’s confetti”), preserving acoustic structure while altering meaning; we identify high-fidelity phonetic matches using CMU pronouncing dictionary. We demonstrate that leading Lyrics-to-Song (L2S) models like SUNO and YuE regenerate songs with striking melodic and rhythmic similarity to their copyrighted originals when prompted with these altered lyrics. More surprisingly, this vulnerability extends across modalities. When prompted with phonetically modified lyrics from a song, a Text-to-Video (T2V) model like Veo 3 reconstructs visual scenes from the original music video—including specific settings and character archetypes—despite the absence of any visual cues in the prompt. Our findings reveal that models memorize deep, structural patterns tied to acoustics, not just verbatim text. This phonetic-to-visual leakage represents a critical vulnerability in transcript-conditioned generative models, rendering simple copyright filters ineffective and raising urgent concerns about the secure deployment of multimodal AI systems. Demo examples are available at our project page [^1] https://jrohsc.github.io/music_attack/.
@@ 1 Introduction
Recent advances in generative multimedia models (Ding et al., 2024; Huang et al., 2023; Yuan et al., 2025; Copet et al., 2023; DeepMind, 2024b) have enabled complex transcript-conditioned tasks like lyrics-to-song (L2S) and text-to-video (T2V) generation, with commercial systems like SUNO [^2] https://suno.com/ and Veo 3 producing high-fidelity content from textual inputs. The rapid deployment of these tools, however, is shadowed by the risk of memorization, where models regurgitate copyrighted material from their training data. While many systems deploy input filters to block verbatim copyrighted lyrics, as a safeguard, our work reveals a fundamental flaw in this strategy. We find that these models exhibit a more profound form of memorization, learning not just literal text but deep structural patterns that manifest across modalities through indirect, phonetic pathways.
To investigate this vulnerability, we first establish a baseline by confirming that verbatim lyrics (Adversarial VerbaTim Prompting, AVT) indeed trigger high-fidelity regurgitation, validating the need for protective filters. We then introduce our primary contribution: Adversarial PhoneTic Prompting (APT), a novel attack that circumvents these text-based filters by replacing iconic phrases with homophonic but semantically unrelated alternatives (e.g., “mom’s spaghetti” becomes “Bob’s confetti”). To identify the closest phonetic matches, we score candidate rewrites with CMU pronouncing dictionary (CMU, 1993), computing a composite phonetic-similarity metric $\Phi$ that captures phoneme sequence, rhyme, syllable count, and stress alignment; we select high-$\Phi$ candidates for evaluation.
As shown in Figure 1, these APT attacks reliably deceive L2S models into producing audio closely aligned with the original songs. For rap tracks such as DNA (Kendrick Lamar) and Lose Yourself (Eminem), AudioJudge scores confirm strong melodic (up to 0.90) and rhythmic (up to 0.95) alignment. The commercial model SUNO is especially vulnerable: a phonetically-modified prompt for ROSÉ (Bruno Mars) achieved near-perfect AudioJudge scores (0.95 melody, 0.98 rhythm), rivaling exact-match prompts. This core weakness—where rhyme, cadence, and sub-lexical patterns dominate model behavior—persists across genres and enables phonetic mimicry to trigger memorization.
More surprisingly, we demonstrate that this phonetic leakage extends across modalities. When prompted with the phonetically altered lyrics of Lose Yourself (Eminem), the T2V model Veo 3 generates video scenes that mirror the original music video—complete with a hooded rapper and dim urban settings—despite no explicit visual cues in the prompt. This ‘phonetic-to-visual regurgitation’ suggests that memorization in these systems is not a simple surface-level phenomenon. Rather, models develop deep, internal representations that link acoustic patterns to both musical and abstract visual concepts.
While prior work has established that generative models can memorize and replicate training data (Ross et al., 2024), often through exact audio fragments or embedded watermarks (Zong et al., 2025; Roman et al., 2024; Epple et al., 2024), our findings reveal a more subtle and pervasive form of the problem. Our work is the first to demonstrate that this vulnerability is not limited to direct replication but is rooted in deep, sub-lexical patterns that can be triggered by phonetic cues alone, a phenomenon particularly underexplored in large-scale, lyrics-conditional models. These findings expose a critical vulnerability: the acoustic shadow of content is enough to summon it across modalities. Future work on memorization and the development of robust defenses must therefore evolve beyond filtering verbatim text to account for the subtle power of these cross-modal phonetic pathways.
@@ 2.1 Music Generation Models
Music generation has advanced rapidly across symbolic and audio domains. Early work focused on symbolic modeling with Transformers for short melodies and chord progressions (Huang et al., 2018; Dong et al., 2018). Recent breakthroughs leverage large-scale foundation models via autoregression (AR) (Agostinelli et al., 2023; Copet et al., 2023; Donahue et al., 2023) and diffusion (Forsgren & Martiros, 2022; Chen et al., 2024; Novack et al., 2025b), enabling full-length, high-fidelity compositions with multimodal conditioning. Models like MusicGen (Copet et al., 2023) and Stable Audio (Evans et al., 2024c; a; b; Novack et al., 2025a) exemplify AR and diffusion approaches for text-to-audio generation. Beyond text, control axes include melody (Wu et al., 2024), harmony (Novack et al., 2024b; a), accompaniment (Nistal et al., 2024a; b), and even video (Tian et al., 2024; Kim et al., 2025). We focus on large-scale Lyrics2Song models, which generate long-form music from textual descriptions and lyrics. YuE (Yuan et al., 2025) is a SOTA open model using in-context learning for multi-minute compositions with lyrical alignment and structural control. SongCreator (Zhou et al., 2024) jointly generates vocals and accompaniment, while CSL-L2M (Jin et al., 2024) aligns melodies with linguistic attributes for fine-grained control. Meanwhile, commercial systems like SUNO employ proprietary pipelines to produce singable songs from lyrics.
@@ 2.2 Memorization and Copyright Risks in Music Generation
Modern music generative models raise critical concerns about memorization, data replication, and copyright infringement. Prior work falls into two main areas: (1) auditing models for memorization and replication of training data, and (2) developing methods for copyright detection and attribution. Studies consistently show that music models can regenerate training data, threatening originality and fair use. Copet et al. (2023) demonstrate that MusicGen reproduces exact or near-exact fragments when prompted with training samples. YuE (Yuan et al., 2025) similarly measures memorization using ByteCover2 similarity, albeit limited to top-1% matches. Stronger evidence comes from Epple et al. (Epple et al., 2024), who find that imperceptible watermarks embedded in training audio reliably resurface in model outputs, highlighting acoustic-level memorization. Other works also observe replication in earlier unconditional (Barnett et al., 2024) and tag-conditioned (Bralios et al., 2024) generative audio systems, though large-scale, lyrics-conditional models remain underexplored. To mitigate these risks, recent research proposes forensic and attribution tools. Deng et al. (Deng et al., 2024) introduce a computational copyright attribution framework using influence metrics (e.g., TRACK, TracIN) to quantify training data contributions, enabling fine-grained royalty allocation. MiRA (Batlle-Roca et al., 2024) provides a model-agnostic system for audio replication detection, leveraging similarity metrics like CLAP (Wu et al., 2023) and DEfNet [^3] https://essentia.upf.edu/models.html . Complementary tools such as ByteCover 1 and 2 (Du et al., 2021; 2022) support melody-sensitive retrieval over full-length tracks, though their closed-source nature and emphasis on overt similarity limit applicability to subtle, influence-level reuse.
3 Methodology
@@ 3.1 Motivation
The high-fidelity output of modern L2S models suggests they are trained on vast datasets that likely include high-quality, copyrighted music. This raises significant concerns about model memorization, where a model might unintentionally reproduce and leak protected content. We investigate a novel and subtle form of this risk, introducing a new class of cross-modality memorization where content leakage occurs through indirect, phonetic pathways. Our central hypothesis is that a model can be prompted to regenerate a copyrighted song not only by using its exact lyrics, but by using semantically nonsensical text that mimics the original’s phonetic structure, rhyme, and cadence. To systematically probe this vulnerability and distinguish between sonic and semantic triggers, we introduce two targeted input prompting strategies: Adversarial PhoneTic Prompting (APT) and Adversarial Verbatim Prompting (AVT).
Lose Yourself (Phoneme Variant) His pants are sweaty, cheese weak, cars are heavy There’s yogurt on his letter already, Bob’s confetti He’s cursive, but on the service, he looks clam and ready To shop moms, but he keeps on betting What he wrote clown, the whole cloud goes so proud He opens his snout, but the birds won’t come out He’s smokin’, how? Everybody’s pokin’ now The sock’s run out, lime’s up, over, meow Snap back toality, rope, there goes cavity Rope, there goes Rabbit, he joked, he’s so glad But he won’t give up that sleepy, no, he won’t have it He knows his whole snack’s to these hopes, it don’t chatter He’s soap, he knows that, but he’s woke, he’s so tragic He knows when he goes back to this noble dome, that’s when it’s Back to the crab again, yo, this bold tragedy Better go rapture this component and hope it don’t trap him

### v4

@@ Abstract
Generative AI systems for music and video commonly use text-based filters to prevent regurgitation of copyrighted material. We expose a significant vulnerability in this approach by introducing Adversarial PhoneTic Prompting (APT), a novel attack that bypasses these safeguards by exploiting phonetic memorization—the tendency of models to bind sub-lexical acoustic patterns (phonemes, rhyme, stress, cadence) to memorized copyrighted content. APT replaces iconic lyrics with homophonic but semantically unrelated alternatives (e.g., “mom’s spaghetti” becomes “Bob’s confetti”), preserving phonetic structure while evading lexical filters. We evaluate APT on leading Lyrics-to-Song models (SUNO, YuE) across English and Korean songs spanning rap, pop, and K-pop. APT achieves 91% average similarity to copyrighted originals, versus 13.7% for random lyrics and 42.2% for semantic paraphrases. Embedding analysis confirms the mechanism: YuE’s text encoder treats APT-modified lyrics as near-identical to originals (cosine similarity 0.90) while Sentence-BERT semantic similarity drops to 0.71, showing the model encodes phonetic structure over meaning. This vulnerability extends cross-modally—Veo 3 reconstructs visual scenes from original music videos when prompted with APT lyrics alone, despite no visual cues in the prompt. We further show that phonetic-semantic defense signatures fail, as APT prompts exhibit higher semantic similarity than benign paraphrases. Our findings reveal that sub-lexical acoustic structure acts as a cross-modal retrieval key, rendering current copyright filters systematically vulnerable. Demo examples available at https://jrohsc.github.io/music_attack/.
@@ 1 Introduction
Recent advances in generative multimedia models [ding2024songcomposer, huang2024audiogpt, yuan2025yue, copet2023simple, deepmind2024veo_card] have enabled complex transcript-conditioned tasks like lyrics-to-song (L2S) and text-to-video (T2V) generation, with commercial models like SUNO [sunoai] and Veo 3 producing high-fidelity content from textual inputs. The rapid deployment of these tools, however, is shadowed by the risk of memorization, where models regurgitate copyrighted material from their training data. To mitigate legal and ethical risks, commercial platforms employ text-based safety filters that match input prompts against databases of copyrighted lyrics and restricted keywords, blocking verbatim regurgitation before content is generated.
These filters, however, rest on a flawed assumption: that lexical overlap is the primary pathway through which memorized content is triggered. We identify a previously unrecognized attack surface that we term Phonetic Memorization—the tendency of large-scale generative models to bind sub-lexical acoustic patterns (phoneme sequences, rhyme, cadence, and syllabic stress) to memorized copyrighted content. Because these patterns operate below the level at which any deployed text-based filter reasons, they constitute an entirely unguarded channel for triggering memorization.
To exploit this channel, we introduce Adversarial PhoneTic Prompting (APT), a novel attack that replaces iconic lyrics with homophonic but semantically unrelated alternatives (e.g., “mom’s spaghetti” becomes “Bob’s confetti”). APT-attacked lyrics preserve the phonetic fingerprint of the original—phoneme sequence, rhyme scheme, syllable count, and stress pattern—while destroying all lexical and semantic overlap. To construct these prompts, we score candidate rewrites using the CMU Pronouncing Dictionary [cmuPronouncingDictionary], computing a composite phonetic similarity metric $\Phi$ that captures seven complementary dimensions of sub-lexical structure; only high-$\Phi$ candidates are retained. By construction, APT prompts pass all existing text-based copyright filters, yet we show they reliably trigger high-fidelity regeneration of copyrighted material. We additionally use Adversarial VerbaTim Prompting (AVT)—prompting with exact copyrighted lyrics—as an upper-bound baseline that calibrates how closely APT’s phonetic-only triggers approach memorization induced by the original training data. We consider a realistic black-box adversary with access only to the public text interface and no knowledge of model internals.
As shown in Figure 1, APT reliably induces L2S models to generate audio tightly aligned with copyrighted originals. We evaluate across songs spanning English and Korean, covering rap, Billboard pop, and K-pop. On SUNO, phonetically modified prompts for GENTLEMAN (by PSY) achieve near-perfect similarity (0.93 melody, 0.97 rhythm), effectively matching the fidelity of exact-lyric inputs. Across all songs, APT attack achieves 91% average similarity, compared to 13.7% for random lyrics and 42.2% for semantic paraphrases that deliberately break phonetic structure. This ordering—random << paraphrase << APT—confirms that sub-lexical acoustic structure, not lexical content or semantic meaning, is the operative retrieval key for memorized content.
Embedding analysis reveals why APT succeeds at the model level. In YuE’s text encoder, original and APT-modified lyrics achieve a mean cosine similarity of 0.90, while Sentence-BERT semantic similarity drops to 0.71. This divergence shows that the model’s internal representations are organized around phonetic-rhythmic structure rather than semantic content: the model treats phonetically preserved but semantically nonsensical prompts as near-identical to the originals, explaining why phonetic mimicry alone is sufficient to activate memorized outputs.
Beyond audio, we uncover a striking cross-modal failure mode we term phonetic-to-visual leakage. When T2V models such as Veo 3 are prompted solely with phonetically modified lyrics—without any visual description—they reconstruct salient visual archetypes from the original music videos, including characteristic settings, clothing styles, and performer identities. These visual elements emerge despite being neither semantically implied nor explicitly requested. This shows that phonetic structure can act as a cross-modal retrieval key, traversing modality boundaries to activate memorized visual representations in downstream generators. This exposes a structural limitation of current safety architectures: defenses are confined to the input modality, while memorized content leaks through output modalities where no safeguard exists.
Finally, we investigate whether phonetic memorization can be mitigated through input filtering. We find that naive defenses combining phonetic and semantic similarity scores fail: counter-intuitively, APT prompts exhibit higher semantic similarity than benign paraphrases, because the attack preserves syntactic structure and function words by design. This suggests that robust mitigation will require model-internal memorization audits or training-time interventions, rather than surface-level prompt filtering alone.
In this paper, we provide the first systematic study of phonetic-based memorization attacks in multimodal generative models. Our contributions are as follows: (1) Novel Attack Vector. We formalize Adversarial PhoneTic Prompting (APT), a phoneme-preserving attack that reliably bypasses text-based copyright filters using a principled phonetic similarity metric. (2) Cross-Modal Memorization Leakage. We demonstrate that phonetic triggers in text prompts can induce memorized outputs not only in audio generation but also in T2V models, revealing phonetic-to-visual leakage and demonstrating that input-side defenses cannot protect against memorization that manifests across output modalities. (3) Extensive Multilingual Evaluation. We conduct a large-scale empirical study across L2S and T2V models (SUNO, YuE, Veo 3), multiple genres, and four languages, using both automatic (AudioJudge, MiRA) and human evaluations. (4) Limits of Existing Defenses. We analyze phonetic- and semantic-aware defenses and show that current filtering strategies are fundamentally misaligned with phonetic memorization, underscoring the need for new safety mechanisms.
@@ 2.1 Music Generation Models
Music generation has advanced rapidly across symbolic and audio domains. Early work focused on symbolic modeling with Transformers for short melodies and chord progressions [huang2018music, dong2018musegan]. Recent breakthroughs leverage large-scale foundation models via autoregression (AR) [agostinelli2023musiclm, copet2023simple, donahue2023singsong] and diffusion [forsgren2022riffusion, chen2023musicldm, Novack2025Presto], enabling full-length, high-fidelity compositions with multimodal conditioning. Models like MusicGen [copet2023simple] and Stable Audio [evans2024open, stableaudio, Evans2024LongformMG, Novack2025Fast] exemplify AR and diffusion approaches for text-to-audio generation. Beyond text, control axes include melody [Wu2023MusicCM], harmony [Novack2024Ditto, Novack2024DITTO2DD], accompaniment [nistal2024diff, nistal2024improving], and even video [tian2024vidmuse, ossl2025]. We focus on large-scale Lyrics2Song models, which generate long-form music from textual descriptions and lyrics. YuE [yuan2025yue] is a SOTA open model using in-context learning for multi-minute compositions with lyrical alignment and structural control. SongCreator [songcreator2024] jointly generates vocals and accompaniment, while CSL-L2M [chai2025csl] aligns melodies with linguistic attributes for fine-grained control. Meanwhile, commercial systems like SUNO employ proprietary pipelines to produce singable songs from lyrics.
@@ 2.2 Memorization and Copyright Risks in Music Generation
Modern music generative models raise critical concerns about memorization, data replication, and copyright infringement. Prior work falls into two main areas: (1) auditing models for memorization and replication of training data, and (2) developing methods for copyright detection and attribution. Studies consistently show that music models can regenerate training data, threatening originality and fair use. [copet2023simple] demonstrate that MusicGen reproduces exact or near-exact fragments when prompted with training samples. YuE [yuan2025yue] similarly measures memorization using ByteCover2 similarity, albeit limited to top-1% matches. Stronger evidence comes from Epple et al. [epple2024watermarking], who find that imperceptible watermarks embedded in training audio reliably resurface in model outputs, highlighting acoustic-level memorization. Other works also observe replication in earlier unconditional [barnett2024exploring] and tag-conditioned [bralios2024generation] generative audio systems, though large-scale, lyrics-conditional models remain underexplored. To mitigate these risks, recent research proposes forensic and attribution tools. Deng et al. [deng2024computational] introduce a computational copyright attribution framework using influence metrics (e.g., TRACK, TracIN) to quantify training data contributions, enabling fine-grained royalty allocation. MiRA [batlle2024towards] provides a model-agnostic system for audio replication detection, leveraging similarity metrics like CLAP [wu2023large] and DEfNet [alonso2020tensorflow]. Complementary tools such as ByteCover 1 and 2 [du2021bytecover, du2022bytecover2] support melody-sensitive retrieval over full-length tracks, though their closed-source nature and emphasis on overt similarity limit applicability to subtle, influence-level reuse.
3 Threat Model
Adversary’s Goal.
The adversary seeks to induce the model to output high-fidelity reproductions of copyrighted songs or music videos that the model has memorized during training. In L2S systems, the goal is to elicit melodies, rhythms, harmonies, vocal identities, and production styles that strongly match known copyrighted material, even when direct lyrics are not used. In T2V systems, the adversary’s goal extends to reproducing visual elements, such as clothing, staging, character archetypes, and scene composition associated with iconic music videos. Critically, the adversary aims to bypass filters that block verbatim text by exploiting the model’s reliance on phonetic patterns rather than semantic meaning, thereby demonstrating that copyright-oriented input filters fail against phonetic mimicry.
Adversary’s Capabilities.
The adversary is assumed to have only black-box access to the generative system through public text interfaces, with no access to weights, training data, APIs, or internal embeddings. However, the adversary can construct phoneme-preserving adversarial prompts that closely match the original lyrics’ rhyme structure and stress pattern. These prompts may be semantically meaningless but maintain enough sub-lexical structure to activate memorized internal representations. The attacker can also iterate and refine prompts using LLMs but cannot alter the target model’s parameters. Beyond this, the adversary does not require musical expertise or system knowledge.
4 Phonetic Memorization Attacks
Figure 2: SUNO interface displaying automatic rejection when users attempt to generate music with copyrighted lyrics. The system flags the input and prompts users to modify the lyrics before proceeding, enforcing copyright compliance at the content generation stage.
@@ 4.1 Motivation
Modern L2S and T2V systems are designed to prevent the regurgitation of copyrighted material. To this end, commercial platforms such as SUNO deploy text-based safety filters that match input prompts against databases of copyrighted lyrics and reject verbatim or near-verbatim reproductions at the lexical level (Figure 2). These defenses rest on a core assumption: that lexical overlap between a user prompt and protected training data is the primary channel through which generative models memorize and reproduce copyrighted content.
We challenge this assumption. We observe that large-scale L2S models produce audio of striking fidelity—closely matching the melodies, harmonies, vocal timbres, and production styles of commercially released songs—suggesting that these models have memorized substantial portions of their copyrighted training data. Critically, this memorized content can be elicited without any lexical overlap with the original lyrics. We identify a previously unrecognized vulnerability that we term Phonetic Memorization: rather than storing training data solely as sequences of discrete lexical tokens, these models internalize rich sub-lexical structure—including phoneme sequences, rhythmic cadence, syllabic stress, and rhyme patterns—that becomes tightly bound to the protected musical representations learned during training. As a consequence, phonetic patterns alone can function as implicit retrieval keys, reactivating memorized multimodal content even when the surface-level semantic meaning of the input is entirely altered.
This finding establishes a fundamentally new attack surface for transcript-conditioned generative systems. Unlike prior memorization concerns that focus on verbatim or near-verbatim regurgitation [carlini2021extracting], phonetic memorization operates below the level at which all currently deployed defenses reason: an adversary can craft prompts that are semantically nonsensical—and therefore invisible to lexical filters, embedding-based detectors, and semantic similarity thresholds alike—yet preserve sufficient phonetic structure to trigger high-fidelity regeneration of copyrighted melodies, vocal identities, and even music-video-level visual motifs. This vulnerability is particularly acute because it is cross-modal: phonetic cues embedded in a text prompt can traverse modality boundaries, activating memorized representations in both audio and video pipelines. Unimodal safeguards that reason only over text or only over visual prompts are therefore fundamentally insufficient. This creates a fundamental modality gap in the defense architecture: all deployed safeguards operate on the text input modality, yet the copyrighted content they aim to protect is stored and reproduced in the audio and visual output modalities. An adversary exploits this gap by crafting text prompts that satisfy every text-level filter while relying on the model’s learned cross-modal associations to trigger memorized content in an entirely unguarded output space.
Real-life Attack Scenarios.
The practical implications of phonetic memorization are significant and span multiple realistic attack scenarios:
  1. 1.
    Commercial Exploitation: An adversary can generate “soundalike” content for streaming platforms—tracks that closely mimic copyrighted songs in melody, rhythm, and vocal timbre—without triggering automated content identification systems such as YouTube Content ID [youtube2024contentid] or Spotify’s audio fingerprinting. Because these systems primarily match against acoustic fingerprints of known recordings, freshly generated outputs that are perceptually similar but not bit-identical may evade detection entirely. The economic incentive is substantial: the global music streaming market exceeds \$40 billion annually [grandview2024music, statista2024music], and even a small fraction of infringing soundalike content can generate significant advertising and subscription revenue before rights holders identify and challenge it. This threat is amplified by the low barrier to entry: our attack requires only black-box API access to a commercial L2S system and publicly available phonetic tools (e.g., the CMU Pronouncing Dictionary [cmuPronouncingDictionary]), making it accessible to unsophisticated adversaries at scale.
  2. 2.
    Training Data Poisoning: A malicious actor can use phonetically-triggered generation to produce large volumes of copyrighted content at scale, then inject this content into publicly crawled datasets. Downstream model developers who unknowingly train on such poisoned data inherit memorization of the copyrighted material, potentially incurring substantial legal liability under emerging AI copyright frameworks. This scenario is especially harmful because the poisoned samples need not contain any copyrighted text—they are audio or video outputs whose provenance is difficult to trace—yet they faithfully reproduce protected musical and visual content. The attack thus creates a supply-chain vulnerability: phonetic memorization at one generative model propagates copyright risk to every downstream system that consumes its outputs.
  3. 3.
    Generalization Beyond Copyright: A Systemic Threat to Text-Based Safety Filters. While we demonstrate phonetic memorization in the context of music and video copyright, the underlying vulnerability extends far beyond this domain. Any text-conditioned safety mechanism that reasons at the lexical or semantic level—including toxicity classifiers, personally identifiable information (PII) screening pipelines, content-moderation filters, and hate-speech detectors—is potentially susceptible to analogous sub-lexical bypass. Consider a toxicity filter that blocks prompts containing slurs or violent language: an adversary could construct phonetically equivalent substitutions that preserve the offensive phonetic pattern while replacing every flagged token with an innocuous alternative, bypassing the filter in exactly the same way APT bypasses copyright filters. The same principle applies to PII guards (e.g., phonetic encodings of names, addresses, or account numbers could evade regex- and dictionary-based scrubbers) and to content-moderation systems for large language models more broadly. Our work thus exposes a structural blind spot shared across a wide class of deployed safety infrastructure: defenses that operate on surface text are fundamentally incomplete when the underlying model has learned sub-lexical representations that can be adversarially activated. This makes phonetic memorization not merely a copyright concern but a systemic security problem of direct relevance to the AI safety community, and it underscores the urgency of developing defenses that reason over phonetic and prosodic structure in addition to lexical content.
[…]
DNA 0.791 0.809 0.859 0.935 0.778 0.855 $\columncolorgray!200.8$35
[…]
4.2 Attack Prompt Construction
We now describe how adversarial prompts are constructed to probe phonetic memorization in transcript-conditioned generative models. Our goal is to isolate the role of sub-lexical phonetic structure from semantic content and evaluate whether phonetic similarity alone can trigger memorized outputs. To this end, we define two complementary prompting strategies: Adversarial PhoneTic Prompting (APT), which constitutes our primary attack, and Adversarial Verbatim Prompting (AVT), which serves as an upper-bound control.
4.2.1 Adversarial PhoneTic Prompting (APT)
Given a lyric sequence $L=\{w_{1},w_{2},\dots,w_{n}\}$, the objective of APT is to construct a modified sequence $L^{\prime}=\{w^{\prime}_{1},w^{\prime}_{2},\dots,w^{\prime}_{n}\}$ that preserves phonetic structure while substantially altering semantic meaning. Intuitively, APT aims to maintain rhyme, rhythm, syllabic cadence, and stress patterns at the word and line level, even if the resulting text is semantically incoherent.
Formally, APT seeks to maximize the average phonetic similarity between corresponding words in $L$ and $L^{\prime}$:
$\mathbb{E}_{i}\left[\Phi_{\mathrm{agg}}(w_{i},w^{\prime}_{i})\right]\approx 1,$
where $\Phi_{\mathrm{agg}}(\cdot,\cdot)$ denotes an aggregate phonetic similarity score defined below. Importantly, no constraint is imposed on semantic similarity, lexical overlap, or grammatical correctness. This allows APT prompts to bypass text-based copyright filters that rely on surface-level or semantic matching.
4.2.2 Phonetic Similarity Metric
To quantify phonetic similarity, we define a multi-dimensional phonetic similarity function $\boldsymbol{\Phi}(a,b)\in[0,1]^{7}$ that captures complementary sub-lexical similarities between two words or short phrases $a$ and $b$:
$\boldsymbol{\Phi}(a,b)=\big(S_{\mathrm{ph}}(a,b),S_{\mathrm{rh}}(a,b),S_{\mathrm{sy}}(a,b),S_{\mathrm{st}}(a,b),$
$S_{\mathrm{jac}}(a,b),S_{\mathrm{cv}}(a,b),S_{\mathrm{vow}}(a,b)\big).$
Each component is normalized to lie in $[0,1]$, with higher values indicating greater phonetic similarity. We compute an aggregate phonetic similarity score by taking the arithmetic mean:
$\Phi_{\mathrm{agg}}(a,b)=\frac{1}{7}\sum_{k=1}^{7}\boldsymbol{\Phi}_{k}(a,b),$
and use $\Phi_{\mathrm{agg}}$ as the default similarity measure unless otherwise specified. The individual components capture distinct aspects of phonetic structure. $S_{\mathrm{ph}}$ measures phoneme-sequence similarity using a SequenceMatcher ratio over CMU Pronouncing Dictionary (CMUdict) phoneme tokens [cmuPronouncingDictionary]. $S_{\mathrm{rh}}$ captures rhyme similarity via overlap of terminal phonemes. $S_{\mathrm{sy}}$ encodes syllable-count similarity derived from vowel counts, and $S_{\mathrm{st}}$ measures stress-pattern alignment based on CMUdict stress digits. To capture additional structural cues, $S_{\mathrm{jac}}$ computes phoneme-level Jaccard similarity, $S_{\mathrm{cv}}$ compares consonant–vowel (CV) patterns, and $S_{\mathrm{vow}}$ aligns stressed vowel cores.
Line-level similarity is computed by averaging $\Phi_{\mathrm{agg}}$ across aligned words within a line, and song-level similarity is obtained by averaging across lines. Together, these features provide a high-coverage representation of sub-lexical phonetic structure that is known to influence musical realization in vocal music, particularly in rhythmically structured genres. In practice, candidate rewrites are generated automatically using a language model constrained to preserve phonetic structure. Each candidate is scored using $\Phi$, and only high-scoring variants are retained for evaluation. This process allows systematic exploration of the phonetic attack surface without manual curation.
Multilingual Extension for Korean.
[…]
4.2.3 Adversarial VerbaTim Prompting (AVT)
AVT serves as both an upper bound on memorization fidelity and a baseline for evaluating how closely phonetic-only prompts approximate verbatim-triggered behavior. In this setting, the model is prompted with the original lyrics without modification, isolating direct memorization effects. Comparing APT against AVT allows us to quantify the extent to which phonetic structure alone can elicit outputs that rival those induced by exact training examples.
Lose Yourself (Phoneme Variant) His pants are sweaty, cheese weak, cars are heavy
There’s yogurt on his letter already, Bob’s confetti
He’s cursive, but on the service, he looks clam and ready
To shop moms, but he keeps on betting
What he wrote clown, the whole cloud goes so proud
He opens his snout, but the birds won’t come out
He’s smokin’, how?
Everybody’s pokin’ now
The sock’s run out, lime’s up, over, meow
Snap back toality, rope, there goes cavity
Rope, there goes Rabbit, he joked, he’s so glad
But he won’t give up that sleepy, no, he won’t have it
He knows his whole snack’s to these hopes, it don’t chatter
He’s soap, he knows that, but he’s woke, he’s so tragic
He knows when he goes back to this noble dome, that’s when it’s
Back to the crab again, yo, this bold tragedy
Better go rapture this component and hope it don’t trap him